# A New Look at the Quantum Measurement Problem

### v2

@@ Abstract
A novel solution to the quantum measurement problem is presented by using a new asymmetric equation that is complementary to the Schrödinger equation. Solved for the hydrogen atom, the new equation describes the temporal and spatial evolution of the wavefunction, and the latter is used to calculate the radial probability density for different measurements. The obtained results are consistent with Born’s position measurement postulates. Specific experimental verification is also proposed.
@@ 1 Introduction
The measurement problem in nonrelativistic quantum mechanics (NRQM) has been an active area of research for almost 100 years. Numerous attempts to solve it (e.g., [1-22]) have been unsuccessful. The proposed ideas are aligned to different interpretations of NRQM, and they range from the objective wavefunction collapse theories, such as the Copenhagen and Ghirardi-Rimini-Weber (GRW) interpretations [2,7-9,13], to the deterministic DeBroglie-Bohm interpretation with its non-local hidden variables [9,13], or to quantum decoherence [15-17], and Everett’s many-words interpretation [3,13]. Penrose [21] proposed a collapse theory that attributes the wavefunction collapse to the influence of gravity. The role of quantum gravity in solving the measurement problem was further explored in [22]; however, this requires to formulate quantum gravity as a fundamental theory of modern physics, which has not yet been accomplished.
@@
According to the Copenhagen interpretation, quantum measurement involves the projection of a previously undetermined quantum state into some measurement eigenstate (e.g., [2,9,13]). This cannot be achieved by the Schrödinger equation (SE) alone. The reason is that the time-evolution of a quantum state $|\Psi>=\hat{U}|\Psi_{o}>$, where $\hat{U}$ is the time-evolution operator for the equation, is continuous and deterministic; however, the projection of $|\Psi>$ into a measurement eigenstate is discontinuous and indeterministic. Thus, von Neumann [2] introduced the reduction postulate, which combines the SE and the collapse of the wavefunction. This leads to ambiguity when a measurement must be switched on and considered as an interaction [9,13]. A possible solution to this problem and its underlying ambiguity is proposed in this paper.
@@
The proposed solution is based on a new asymmetric equation (NAE) [23], whose existence is guaranteed by the irreducible representations (irreps) of the Galilean group [24-32]. The NAE is complementary to the SE [33,34] that also originates from the same irreps [31,32]. The main advantage of the NAE, as compared to the SE, is that it allows measurements to be directly accounted for. The effects of the measurements on the evolution of the wavefunction are investigated by solving the new equation for the hydrogen atom. The obtained solutions represent the temporal and spatial evolution of the wavefunction, and they are used to compute the radial probability densities of finding the electron after the measurement process is completed.
@@
2 Basic equations from the Galilean group
@@
The irreducible representations (irreps) of the Galilean group, $\mathcal{G}$, are well-known and classified [24-30]. A scalar function $\phi(t,\mathbf{x})$ transforms as one of the irreps of $\mathcal{G}$ if the following eigenvalue equations $i\partial_{t}\phi=\omega\phi$ and $-i\nabla\phi=\mathbf{k}\phi$ are satisfied, where $\partial_{t}=\partial/\partial t$, and $\omega$ and $\mathbf{k}$ are the eigenvalues that label the irreps [31,32]. Using the eigenvalue equations only two second-order partial differential equations that are asymmetric in space and time derivatives can be derived. The coefficients in these equations are expressed as ratios of the labels of the irreps. If these labels are identified as wave frequencies and wavenumbers, the resulting asymmetric equations describe propagation of classical waves [35].
However, if the coefficients are determined with the de Broglie relationship (e.g., [33,34]), then, one of these asymmetric equations becomes the Schrödinger equation
@@
There are several important differences between the SE and NAE. The most prominent one is the difference in time derivatives. By being first-order in time, the time solutions to the SE are also the solutions to the continuity equation for probability, which is also first-order in time. This means that the SE is invariant with respect to unitary transformations, and that NRQM based on the SE is a unitary theory. On the other hand, the NAE being second-order in time may have solutions that do not satisfy the continuity equation and, as a result, the NAE may be used to describe some non-unitary processes in NRQM. Thus, the SE (unitary) $\rightarrow$ the NAE (non-unitary), with the latter describing processes as the measurement problem or the quantum jumps representing absorption or emission of radiation by atoms (e.g., [33,34]); only the former is considered in this paper.
@@ 3 Governing equation and measurements
The labels of the irreps, $\omega$ and $\mathbf{k}$, which appear explicitly in the NAE can be used to specify interactions between a quantum system and an apparatus that performs measurements on this system. Denoting $\omega=\omega_{ap}$, $E_{ap}=\hbar\omega_{ap}$, and ${\mathbf{k}}={\mathbf{k_{ap}}}$, with ${\mathbf{k_{ap}}}={\hat{k}(1/\lambda_{ap}})$, where the parameters $\omega_{ap}$ and $\lambda_{ap}$ may represent frequency and wavelength of electromagnetic (EM) waves used in the measurements, respectively, or be determined by temporal and spatial scales associated with the measuring process. Then, the eigenvalue equations for the energy, $\hat{E}=i\hbar\partial_{t}$, and momentum, $\mathbf{\hat{P}}=-i\hbar\nabla$, operators acting on the wavefunction $\phi_{A}$ are given by $\hat{E}\phi_{A}=\sqrt{E_{ap}E}\ \phi_{A}$ and $(\hbar\mathbf{k_{ap}}\cdot\mathbf{\hat{P}})\phi_{A}=(\mathbf{p_{ap}}\cdot\mathbf{\hat{P}})\phi_{A}=p_{ap}p\ \phi_{A}$, with $E=\hbar\omega$, $p=\hbar k$, and $p_{ap}=\hbar k_{ap}$ [37].
@@
The Schrödinger equation with the Coulomb potential, $V(r)=-e^{2}/(4\pi\epsilon_{o}r)$, describes a hydrogen atom by giving the solutions for the wavefunction in terms of Laguerre’s polynomials (the radial-dependence) and spherical harmonics (the angular-dependence), and predicts correctly its quantum energy levels given by $E_{n}=-(m/2n^{2}\hbar^{2})(e^{2}/4\pi\epsilon_{o})^{2}$. The results given by the SE descibe the electron’s unitary dynamics and they are valid prior to making any measurement.
@@
However, the probability densities resulting from the measurements, and calculated with the NAE, do not account for the previous electron’s unitary dynamics. To take this into consideration, the probabilities resulting from the measurements are multipled by the original probablility density obtained for the electron from the SE. The resulting probabliltes demonstrate that the original state of the electron is projected into the observable eigenstate as suggested by Born’s measurement postulates (e.g., [33.34]).
@@ 4 Time-dependent solutions and their physical meaning
where $E_{ap}=\hbar\omega_{ap}$ and the Bohr radius $a_{o}=(4\pi\epsilon_{o}\hbar^{2})/me^{2}=\hbar/(mc\alpha)$, with $\alpha$ being the fine structure constant. It must be noted that both $\omega_{ap}$ and $k_{ap}$ are specified by interactions between a measuring apparatus and a quantum system.
@@
In general, it is expected that the time required for the wavefunction collapse is shorter when $m$ increases (e.g., [9,13,33]). The arguments of the periodic and exponential parts of the solution ${\cal{R}\it e}[\chi(t)]$ are the same, and they depend explictly on the mass of electron $m$. Thus, with $a_{o}$ being a function of $m$, the arguments depend on $\sqrt{m}$, whose larger value causes the exponential function to decrease faster, and shortens the time of the wavefunction collapse.
@@ 5 Time-indepedent solutions and probability densities
$I_{n}(\beta_{ap})=\int_{0}^{\infty}r_{a}^{4\beta_{ap}+2}\ e^{-2(\beta_{ap}/n^{2})r_{a}}=\frac{1}{(2\beta_{ap})^{4\beta_{ap}+3}}\ \Gamma({4\beta_{ap}+3})\ ,$ (14)
which is valid if ${\cal{R}\it e}(2\beta_{ap})>0$ and ${\cal{R}\it e}(4\beta_{ap}+2)>-1.$
@@
[…]
@@
The above procedure and calculations can also be performed for the principal quantum numbers $n=2$, $3$, $4$, and so on, which will give different values of $r_{max}$, and different probability densities; such calculations will be presented elsewhere.

### v3

# A Solution to the Quantum Measurement Problem

@@ Abstract
A novel solution to the quantum measurement problem is presented by using a new asymmetric equation that is complementary to the Schrödinger equation. Solved for the hydrogen atom, the new equation describes the temporal and spatial evolution of the wavefunction, and the latter is used to calculate the radial probability density for different measurements. The obtained results show that Born’s position measurement postulates naturally emerge from the theory and its first principles. Experimental verification of the theory and its predictions is also proposed.
@@ 1 Introduction
The measurement problem in nonrelativistic quantum mechanics (NRQM) has been an active area of research for almost 100 years. Numerous attempts to solve it (e.g., [1-29]) have been unsuccessful. The proposed ideas are aligned to different interpretations of NRQM, and they range from the objective wavefunction collapse theories such as the Ghirardi-Rimini-Weber (GRW) model [8,9] and its extension known as the continuous spontaneous localization (CSL) model [10,11] that are based on the Copenhagen interpretation [2-7]], to the deterministic DeBroglie-Bohm interpretation with its non-local hidden variables [7,15], or to quantum decoherence [17-19], and Everett’s many-words interpretation [3,15]. Penrose [23] proposed a collapse theory that attributes the wavefunction collapse to the influence of gravity. The role of quantum gravity in solving the measurement problem was further explored in [24]; however, this requires to formulate quantum gravity as a fundamental theory of modern physics, which has not yet been accomplished.
@@
According to the Copenhagen interpretation, quantum measurements involve the projection of a previously undetermined quantum state into an eigenstate that corresponds to the measurements (e.g., [2,7,15]). This cannot be achieved by the Schrödinger equation (SE) alone. The reason is that the time-evolution of a quantum state $|\Psi>=\hat{U}|\Psi_{o}>$, where $\hat{U}$ is the time-evolution operator for the equation, is continuous and deterministic; however, the projection of $|\Psi>$ into an eigenstate corresponding to the measurements is discontinuous and indeterministic. Heisenberg [25] argued that an energy transfer between a detector and an observed quantum system disrupts the smooth evolution of the wavefunction through quantum jumps (see [26] for a possible resolution]).
On the other hand, von Neumann [2] tried to solve the measurement problem by introducing the reduction postulate, which combines the SE and the collapse of the wavefunction. The reduction postulate remains controversial [27], and some arguments were made to demonstrate that the postulate is incorrect (e.g., [28]) because it leads to ambiguity when a measurement must be switched on and considered as an interaction [29]. A solution to the quantum measurement problem is proposed in this paper.
A quantum system interacting with its surroundings is called an open quantum system, and a significant progress in studying such systems has been made during the last few decades, with the Theory of Open Quantum Systems being established (e.g., [30]). Specifically, theoretical and experimental research on open quantum systems by using digital quantum simulators has been performed and general algorithms for such digital simulations have been investigated and developed [31-34]. However, their practical experimental implementation poses some challenges that still remain unsolved [33,34]. The approach described in this paper is very different from the open system approach, and it has several advantages when compared to the latter as the presented theory and its results demonstrate.
@@
A solution of the quantum measurement problem proposed in this paper is based on a new asymmetric equation (NAE) [35], whose existence is guaranteed by the irreducible representations (irreps) of the Galilean group of the metric [36-38]. The NAE is complementary to the SE that also originates from the same irreps [39,40]. The main advantage of the NAE, as compared to the SE, is that it allows measurements to be directly accounted for. The effects of the measurements on the evolution of the wavefunction are investigated by solving the new equation for the hydrogen atom. The obtained solutions represent the temporal and spatial evolution of the wavefunction, and they are used to compute the radial probability densities of finding the electron after the measurement process is completed.
@@
2 Basic equations from the Galilean group of the metrics
Galilean space and time are separated and their metrics are given by $ds_{1}^{2}\ =\ dx^{2}+dy^{2}+dz^{2}$ and $ds_{2}^{2}\ =\ dt^{2}$, where $x$, $y$ and $z$ are spatial coordinates and $t$ is time. The metrics remain invariant with respect to all the transformations that form the Galilean group of the metric, ${\mathcal{G}}$, whose structure can be written as ${\mathcal{G}}\ =S(4)\otimes_{s}H(6)$, where $S(4)=T(1)\otimes R(3)$, with $T(1)$ and $R(3)$ being subgroups of translation in time and rotations, respectively. In addition, $H(6)=T(3)\otimes B(3)$, where $T(3)$ and $B(3)$ are subgroups of translations in 3D space and Galilean boosts, respectively [36-38].
The Galilean group of the metrics, ${\mathcal{G}}$, is a ten-parameter Lie group and $H(6)$ is its invariant Abelian subgroup. The irreducible representations (irreps) of $H(6)$ are one-dimensional and they are well-knonw [39,40], and they provide labels for all the irreps of the entire group ${\mathcal{G}}$. Classification of these irreps was done by Bargmann [36], who demonstrated that the scalar and spinor irreps are physical, but the vector and tensor irreps are not because they do not allow for the elementary particle localization. As a result, in NRQM only scalar and spinor wavefunctions are allowed, and their temporal and spatial evolution is described by the Schrödinger [41,42] and Lévy-Leblond [37,38] equations, respectively. In this paper, only the scalar irreps are considered.
The transformations for space translations represented by $T(3)$, and Galilean boosts represented by $B(3)$, may be considered separately and written as: $\hat{T}_{\mathbf{a}}\phi(\mathbf{r},t)\ \equiv\ \phi(\mathbf{r}+\mathbf{a},t)\ =e^{i\mathbf{k}\cdot\mathbf{a}}\phi(\mathbf{r},t)$ and $\hat{B}_{\mathbf{v}}\phi(\mathbf{r},t)\ \equiv\ \phi(\mathbf{r}+\mathbf{v},t)\ =\ e^{i\mathbf{k}\cdot\mathbf{v}t}\phi(\mathbf{r},t)$, where $\hat{T}_{\mathbf{a}}$ and $\hat{B}_{\mathbf{v}}$ are operators of space translations and boosts, respectively. Moreover, $\phi(\mathbf{r},t)$ is a scalar wavefunction, $\mathbf{a}$ represents a translation in space, $\mathbf{v}$ is the velocity of Galilean boosts, and the real vector $\mathbf{k}$ labels the one-dimensional irreps of $H(6)$.
Expanding $\phi(\mathbf{r}+\mathbf{a},t)$ and $\phi(\mathbf{r}+\mathbf{v}t,t)$ in Taylor series and comparing the results to the above transformations, one obtains the following eigenvalue equation [37,38]
$-i\nabla\phi(\mathbf{}r,t)\ =\ \mathbf{k}\phi(\mathbf{}r,t)\ ,$ (1)
which has the same form for the boosts and translations in space, and it is the necessary condition that the wavefunction $\phi(\mathbf{}r,t)$ transforms as one of the irreps of ${\mathcal{G}}$.
However, $S(4)$ is not an invariant subgroup of ${\mathcal{G}}$, which means that the above procedure cannot be applied to it; the fact that $T(1)$ is an invariant subgroup of $S(4)$ does not help because $T(1)$ is not the ’little group’ of ${\mathcal{G}}$ [39]. The proposed solution [39] requires that $\phi_{\omega}(\mathbf{r},t)=\eta(\mathbf{r},t)\phi(\mathbf{r},t)$, where $\eta(\mathbf{r},t)$ is a smooth and differentiable function to be determined, and $\hat{E}=i\partial_{t}=i\partial/\partial t$ is the generator of translation; then, $i\partial_{t}\phi_{\omega}(\mathbf{r},t)\ =\ \omega\phi_{\omega}(\mathbf{r},t)$. The results presented in [39] demonstrated that $\eta(\mathbf{r},t)=\eta^{\prime}(\mathbf{r}^{\prime},t^{\prime})=1$ in order to be in agreement with the Galilean Principle of Relativity. Thus, the final temporal eigenvalue equation [39] can be written as
$i\partial_{t}\phi(\mathbf{r},t)\ =\ \omega\phi(\mathbf{r},t)\ ,$ (2)
and this equation supplements Eq. (1). The eigenvalue equations given by Eqs (1) and (2) guarantee that the wavefunction $\phi(\mathbf{r},t)$ transforms as one of the irreps of ${\mathcal{G}}$, and that these equations can be used to derive dynamical equations that are consistent with the Galilean group of the metric ${\mathcal{G}}$.
There is another approach to find dynamical equations by using the Casimir operator of ${\mathcal{G}}$, which is well-known (e.g., [39]). However, this operator does not connect Galilean space and time, and therefore, it cannot be used to obtain any dynamical equation. The problem is resolved by extending ${\mathcal{G}}$, so that the connection is allowed. This results in the so-called extended Galilean group, whose structure is $\mathcal{G}_{e}\ =[R(1)\otimes B(3)]\otimes_{s}[T(3+1)\otimes U(1)]$, where $T(3+1)$ is an invariant subgroup of $\mathcal{G}_{e}$ of translations in space and time, and $U(1)$ is a one-parameter unitary group (e.g., [37,38]); note that the structure of this group is similar to the Poincaré group (e.g., [43-46]), and that $\mathcal{G}_{e}$ has three Casimir operators. If a scalar wavefunction is considered, then, one of these Casimir operators gives directly the Schrödinger equation [39], but this requires a prior knowledge of $\mathcal{G}_{e}$, which was introduced only after the Schrödinger equation was already established and used in NRQM. Moreover, it must be also noted that $\mathcal{G}_{e}$ is not the group of the Galilean metrics. Therefore, the derivation of the Schrödinger equation from this Casimir operator may not be consistent with the Principle of Galilean Relativity.
@@ 2 Basic equations from the Galilean group of the metrics
As a result, the eigenvalue equations given by Eqs (1) and (2) are used here to obtain two second-order partial differential equations that are asymmetric in space and time derivatives. The coefficients in these equations are expressed as ratios of the eigenvalues, which label the irreps of $\mathcal{G}_{e}$. If these labels are identified as wave frequencies and wavenumbers, the resulting asymmetric equations describe propagation of classical waves [47]. However, the coefficients can be determined by using the de Broglie relationship (e.g., [41,42]), then, one of the resulting asymmetric equations is the Schrödinger equation
@@
There are several important differences between the SE and NAE. The most prominent one is the difference in time derivatives. By being first-order in time, the time solutions to the SE are also the solutions to the continuity equation for probability, which is also first-order in time. This means that the SE is invariant with respect to unitary transformations, and that NRQM based on the SE is a unitary theory. On the other hand, the NAE being second-order in time may have solutions that do not satisfy the continuity equation and, as a result, the NAE may be used to describe some non-unitary processes in NRQM. Thus, the SE (unitary) $\rightarrow$ the NAE (non-unitary), with the latter describing processes such as the quantum measurement problem investigated in this paper, or quantum jumps that represent absorption and emission of electromagnetic (EM) radiation by atoms, and were recently studied in [47] by using the NAE.
@@ 3 Governing equation and measurements
The parameters $\omega$ and $\mathbf{k}$ that appear explicitly in the NAE (see Eq. 4) are the eigenvalues of the Hermitian operators in the eigenvalue equations given by Eqs (1) and (2), which means that they are real and label the irreps of the group. Thus, they can also be associated with a measuring apparatus, specifically with the EM radiation that this apparatus uses to probe a quantum system. For the theory developed in this paper, it is sufficient to prescribe the frequency of EM radiation $\omega=\omega_{ap}$, which gives $E_{ap}=\hbar\omega_{ap}$, and calculate the resulting wavevector ${\mathbf{k}}={\mathbf{k_{ap}}}$, with ${\mathbf{k_{ap}}}={\hat{k}(1/\lambda_{ap}})$, where $\lambda_{ap}$ is the wavelength of EM used by the apparatus. In general, $\omega$ and $\mathbf{k}$ may also be determined by temporal and spatial scales associated with the measuring process.
Then, the eigenvalue equations for the energy, $\hat{E}=i\hbar\partial_{t}$, and momentum, $\mathbf{\hat{P}}=-i\hbar\nabla$, operators acting on the wavefunction $\phi_{A}$ are given by $\hat{E}\phi_{A}=\sqrt{E_{ap}E}\ \phi_{A}$ and $(\hbar\mathbf{k_{ap}}\cdot\mathbf{\hat{P}})\phi_{A}=(\mathbf{p_{ap}}\cdot\mathbf{\hat{P}})\phi_{A}=p_{ap}p\ \phi_{A}$, with $E=\hbar\omega$, $p=\hbar k$, and $p_{ap}=\hbar k_{ap}$ [47-49].
@@
The Schrödinger equation with the Coulomb potential, $V(r)=-e^{2}/(4\pi\epsilon_{o}r)$, describes a hydrogen atom by giving the solutions for the wavefunction in terms of Laguerre’s polynomials (the radial-dependence) and spherical harmonics (the angular-dependence), and predicts correctly its quantum energy levels given by $E_{n}=-(m/2n^{2}\hbar^{2})(e^{2}/4\pi\epsilon_{o})^{2}$. The results given by the SE descibe the electron’s unitary evolution, and they are valid prior to making any measurement.
@@
However, the probability densities resulting from the measurements, and calculated with the NAE, do not account for the previous electron’s unitary dynamics. To take this into consideration, the probabilities resulting from the measurements are multipled by the original probablility density obtained for the electron from the SE. The resulting probabliltes demonstrate that the original state of the electron is projected into the observable eigenstate, as suggested by Born [50,51]; this is now known as Born’s position measurement postulates or Born’s rules.
Born’s rules suggest that when a quantum particle interacts with a measuring apparatus, then it gets confined into a measuring eigenstate that corresponds to a well-defined position (e.g., [41,42]). In other words, the rules connect NRQM and its theoretical framework to experiments because they allow calculating probablilities from quantum mechanical amplitudes. Thus, they become a crucial link between abstract NRQM and the real world of experience. However, no first principle derivations of the rules exist so far, and they are considered to be Born’s intelligent guess (e.g., 52,53]). In this paper, Born’s rules are described by the NAE, which makes them to originate from the same basic principles as the NAE does, namely, from the irreps of the Galilean group of the metric and the Principles of Galilean Relativity.
@@ 4 Time-dependent solutions and their physical meaning
where $E_{ap}=\hbar\omega_{ap}$, $\mathbf{k_{ap}}=k_{ap}\mathbf{\hat{k}}_{ap}$, $\mathbf{r}=r\mathbf{\hat{r}}$, and the Bohr radius $a_{o}=(4\pi\epsilon_{o}\hbar^{2})/me^{2}=\hbar/(mc\alpha)$, with $\alpha$ being the fine structure constant. It must be noted that both $\omega_{ap}$ and $k_{ap}$ are specified by interactions between a measuring apparatus and a quantum system.
@@
In general, it is expected that the time required for the wavefunction collapse is shorter when $m$ increases (e.g., [7,15,41,42]). The arguments of the periodic and exponential parts of the solution ${\cal{R}\it e}[\chi(t)]$ are the same, and they depend explictly on the mass of electron $m$. Thus, with $a_{o}$ being a function of $m$, the arguments depend on $\sqrt{m}$, whose larger value causes the exponential function to decrease faster, and shortens the time of the wavefunction collapse.
@@ 5 Time-indepedent solutions and probability densities
$I_{n}(\beta_{ap})=\int_{0}^{\infty}r_{a}^{4\beta_{ap}+2}\ e^{-2(\beta_{ap}/n^{2})r_{a}}=\frac{1}{(2\beta_{ap}/n^{2})^{4\beta_{ap}+3}}\ \Gamma({4\beta_{ap}+3})\ ,$ (16)
which is valid if ${\cal{R}\it e}(2\beta_{ap})>0$ and ${\cal{R}\it e}(4\beta_{ap}+2)>-1$; note that both conditions are obeyed in the theory presented in this paper.
@@
Using Eqs (15) and (16), the radial probability density ${\cal{P}}_{1}(r_{a})$ is calculated for $n=1$ and for different values of $\beta_{ap}$, which varies from $50$ to $9000$. This range of $\beta_{ap}$ corresponds to $\lambda_{ap}=50\ a_{o}$ and $100\ a_{o}$ (X-rays), $\lambda_{ap}=200\ a_{o}$ and $800\ a_{o}$ (ultraviolet), and $\lambda_{ap}=8000\ a_{o}$ and $9000\ a_{o}$ (visible); this range can be extended to the other parts of the EM spectrum. Since the selected values of $\beta_{ap}$ are positive integers, the function $\Gamma({4\beta_{ap}+3})=({4\beta_{ap}+2})!$ in Eq. (16). The obtained results are presented in Fig. 1 showing also the radial probability density, ${\cal{P}}_{1s}(r_{a})=4r_{a}^{2}\ e^{-2r_{a}}$, for the hydrogen $1s$ orbital plotted for comparison.
@@
The theory and results presented in this paper descibe the process known as quantum decoherence, which occurs when a quantum system interacts with its environment. As a result of this interaction, the information that the system contains is admixed up with its environment. More specifically, the system loses its ability to show superposition and interference effects, and its physical behavior resembles a classical system. This can be seen in Fig. 2, which shows that the particle is confined to a very narrow eigenstate after measurements are done. Moreover, according to Eq. (17), the longer the wavelength of EM radiation absorbed by the system, the more prominent its classical-like behavior is observed.
@@
Since the presented results depend directly on the principal quantum number $n$ (see Eqs 15 and 16), similar calculations can be performed for $n=2$, $3$, $4$; however, they will be presented elsewhere.